OpenStudy (anonymous):

3a * 9c =27ac A. identity property B. reciprocal property of multiplication C. commutative property of multiplication D. zero property of multiplication

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

I think reciprocal one is like a * b = a * b zero one is a * 0 = 0 you are definitely multiplying something, so which do you think it can be?

OpenStudy (anonymous):

D zero property of multiplication

jigglypuff314 (jigglypuff314):

.... is there a zero anywhere in the equation they gave you?

OpenStudy (anonymous):

thank you so much

OpenStudy (fakereality):

Dang. I'm late again! D:<